5. Use your closest value to calculate a value of the universal gravitational constant G, where the mass of the Earth Me 5.983 X 1024 kg and its radius is r=6.378 X 106 m. ME em. 6.37 10 W UESTIONS Which results gave the more accurate readings? Why is this so?
Answer #1

g = GM/r^2

where g = acceleration due to gravity

G = gravitational constant

M = mass of earth

R = raidus of earth

9.8 = G*5.983*10^24 / (6.378*10^6)^2

G = 6.663*10^-11 m3⋅g1⋅s2